OpenStudy (anonymous):

sum of a and b is a+b, and the product of m and n=m+n, two less than k is=2-k, s divided by n is s /n

myininaya (myininaya):

product means multiply

OpenStudy (anonymous):

so answer is mn

myininaya (myininaya):

and also you wrote k less than 2 instead of 2 less than k

OpenStudy (anonymous):

my bad and 2-k=k-2 or 2-k

myininaya (myininaya):

2-k isn't k-2

myininaya (myininaya):

m less than n means n-m

myininaya (myininaya):

so two less than k means k-2

OpenStudy (anonymous):

sum of a and b is a+b right

myininaya (myininaya):

the 2 things I wrote about are the only things I see wrong

OpenStudy (anonymous):

sum of a and b is a+b right

myininaya (myininaya):

yes as I said the two things I pointed out are the only things I see wrong

myininaya (myininaya):

so that means s divided by n =s/n is also correct

OpenStudy (anonymous):

h increased by 10 is h+10

OpenStudy (anonymous):

quotient of x and 4 is